# Topological transitions in the presence of quenched uncorrelated disorder

### v1

@@ Abstract
We address issues related to the presence of defects at topological transitions, in particular when defects are modeled in terms of further variables associated with a quenched disorder, corresponding to the limit in which the defect dynamics is very slow. As a paradigmatic model, we consider the three-dimensional lattice ${\mathbb{Z}}_{2}$ gauge model in the presence of quenched uncorrelated disorder associated with the plaquettes of the lattice, whose topological transitions are characterized by the absence of a local order parameter. We study the critical behaviors in the presence of weak disorder. We show that they belong to a new topological universality class, different from that of the lattice ${\mathbb{Z}}_{2}$ gauge models without disorder, in agreement with the Harris criterium for the relevance of uncorrelated quenched disorder when the pure system undergoes a continuous transition with positive specific-heat critical exponent.
@@ I Introduction
Statistical systems with quenched disorder, modeling the presence of defects subject to a slow dynamics, generally develops peculiar critical phenomena associated with new universality classes with respect to pure systems, such as transitions to glassy phases. Phase transitions and critical behaviors in disordered systems have been mostly studied within lattice spin models with various realizations of local disorder, in particular with types of disorder preserving the symmetry of the pure system, such as random-bond and random-site disorder (random-field disorder instead generally breaks the global symmetry), and spatially uncorrelated probability distributions of the quenched variables, see, e.g., Refs. Ma-book ; Nishimori-book ; Harris-74 ; EA-75 ; Nishimori-81 ; HPPV-07 ; HPPV-07-rdi ; HPPV-07-mc ; PC-99 ; Betal-00 ; KM-02 ; KR-03 ; KKY-06 ; Jorg-06 ; HPV-08 ; BFMMMY-11 ; Janus-13 ; LPP-16 ; Granato-04 ; MHT-04 ; PR-05 ; LY-07 ; VK-09 ; FMPTY-09 ; BFMS-14 ; OUK-20 ; AV-11 ; Asp-etal-16 ; Hartmann-99 ; CPV-11 ; PHP-06 ; LH-88 ; HPPV-08 ; PPV-09 ; FMMPR-16 ; CP-19 ; Nishimori-24 ; Vojta-19 .
@@
While a large amount of disorder may lead to glassy phases and glassy critical behaviors in three-dimensional (3D) spin systems, ferromagnetic transitions may still occur for sufficiently weak disorder. However, even though the quenched disorder does not change the symmetry-breaking pattern, the universality class of the critical behavior may differ from that of the ferromagnetic transition of the pure system. According to the Harris criterium Harris-74 , which applies to a spatially-uncorrelated quenched disorder that is effectively coupled to the energy density, ferromagnetic transitions in the presence of weak disorder may develop a new universality class when the specific-heat exponent $\alpha$ of the pure transition is positive, as it happens in 3D random-bond and random-site Ising systems Ma-book ; Nishimori-book ; HPPV-07 ; HPPV-07-rdi . On the other hand, if the pure-system continuous transition has a negative specific-heat exponent, then its universality class turns out to be stable with respect to weak disorder, as it happens in 3D random-bond O($N$) vector models for any $N\geq 2$, such as the XY and Heisenberg spin models PV-02 , corresponding to $N=2$ and $N=3$ respectively. Moreover, with increasing disorder, spin systems may develop peculiar multicritical behaviors along the so-called Nishimori line Nishimori-book ; Nishimori-81 ; LH-88 ; HPPV-07-mc ; HPPV-08 ; AV-11 , and then transitions to glassy phases, see, e.g., Refs. KKY-06 ; HPV-08 ; BFMMMY-11 ; Janus-13 ; LPP-16 ; Granato-04 ; MHT-04 ; PR-05 ; LY-07 ; VK-09 ; FMPTY-09 ; BFMS-14 ; OUK-20 ; AV-11 ; Asp-etal-16 ; CPV-11 .
@@
The effects of quenched disorder have been much less investigated at topological transitions, such as those occurring in systems with gauge symmetries, without a local order parameter driving the transition Wegner-71 ; Kogut-79 ; Sachdev-19 ; BPV-25 . A paradigmatic model for systems undergoing finite-temperature topological transitions is provided by the 3D lattice ${\mathbb{Z}}_{2}$ gauge model Wegner-71 ; Kogut-79 , whose confinement-deconfinement transition is not driven by a local order parameter but by the nonlocal order parameter associated with the area/perimeter law of the Wilson loops Wegner-71 ; Kogut-79 . 3D lattice ${\mathbb{Z}}_{2}$ gauge models have been also considered in the presence of local spatially-uncorrelated quenched disorder associated with each plaquette of the lattice DKLP-02 ; WHP-03 (denoted by RPGM in the following), i.e., with randomly chosen plaquettes having couplings with the wrong (negative instead of positive) sign. The RPGM was originally introduced in relation to the theory of the quantum error correction. However, it also provides a paradigmatic example of systems undergoing topological transitions in the presence of disorder. Some results on its temperature-disorder phase diagram were reported in Refs. WHP-03 ; OAIM-04 , but the critical behavior at the transition lines has never been analyzed. Therefore the characterization of the universality classes of its topological transitions in the presence of quenched disorder is still an open issue.
@@
In this paper we investigate the effects of disorder at topological transitions, focusing on the paradigmatic 3D RPGM. We recall that the pure 3D lattice ${\mathbb{Z}}_{2}$ gauge model is related to the standard 3D Ising model by a duality relation Wegner-71 ; Savit-80 , thus they share the same length-scaling critical exponent $\nu_{\cal I}\approx 0.630$ PV-02 , which implies that the specific-heat exponent is positive, $\alpha_{\cal I}\approx 0.110$. We present a numerical analysis of the critical behaviors of the 3D lattice ${\mathbb{Z}}_{2}$ gauge model for weak disorder, which is made difficult by the fact that its topological transitions do not have gauge-invariant local order parameters. We thus study the critical behaviors of this system by focusing on the scaling behavior of the gauge-invariant energy cumulants. The numerical results show that the quenched disorder destabilizes the pure-system topological transition, confirming the Harris criterium. Indeed, we observe critical behaviors belonging to another universality class with $\nu=0.82(2)$, which is significantly larger that the Ising length-scale exponent $\nu_{\cal I}\approx 0.630$ in the absence of disorder (and also any universality class of Ising systems with quenched disorder). In the final part of the paper we also briefly discuss generalizations of the RPGM, considering generic 3D lattice ${\mathbb{Z}}_{N}$ gauge models with weak disorder.
@@ III Numerical methods and results
$B_{k}(K,L)\approx L^{k/\nu-3}{\cal B}_{k}(X)+b_{k},$ (11)
@@
where the constants $b_{k}$ represent the analytic background PV-02 ; BPV-24-coH , and we neglect $O(L^{-\omega})$ scaling corrections to the leading scaling behavior [the exponent $\omega>0$ is generally associated with the leading irrelevant renormalization group (RG) perturbation at the fixed point PV-02 , for example $\omega\approx 0.8$ at the standard Ising transition, i.e., for $q=0$]. The scaling functions ${\cal B}_{k}(X)$ are universal apart from a multiplicative factor and a normalization of the argument, however they generally depend on the boundary conditions adopted.
@@
If $\omega<3/\nu-3$ then $\kappa=\omega$ and the $O(L^{-\kappa})$ term parametrizes the correction due to the leading irrelevant perturbation [not explicitely written in Eq. (11)]. If instead $\omega>3/\nu-3$ then $\kappa=3/\nu-3$ and the $O(L^{-\kappa})$ term takes into account the analytic background in Eq. (11), and in this case $n_{s}=0$ should be used.
@@
instead of Eq. (12). The final result of this analysis are the estimates
@@ IV Conclusions
We have investigated the effects of quenched disorder at topological transitions, focusing on the 3D ${\mathbb{Z}}_{2}$ gauge model in the presence of an uncorrelated quenched disorder associated with the plaquettes of the lattice (RPGM). This is paradigmatic model for systems undergoing finite-temperature topological transitions Wegner-71 ; Kogut-79 ; Sachdev-19 ; BPV-25 , indeed its confinement-deconfinement transitions are not driven by any local order parameter. We report a numerical analysis of their critical behaviors for weak disorder, and for this purpose we focus on the scaling behavior of the gauge-invariant energy cumulants defined in Eq. (6). The numerical investigation has been carried out using the binomial distribution for disorder variables, but it is natural to guess that the universality class of the results is independent of the disorder distribution, as it happens for the Ising model with quenched disorder, see, e.g., Ref. HPPV-07-rdi .
@@
${1\over V}[\langle\mathcal{H}\rangle]_{w}=3-6q$ (17)
holds along the Nishimori line (5) (see also analogous behaviors Nishimori-book ; Nishimori-81 for the random-bond Ising model). Therefore, the scaling part of the cumulants vanishes along the Nishimori. We also note that the critical behaviors along the deconfinement-confinement transition line below the Nishimori point, see Fig. 1, is expected to belong to a different universality class, as it occurs for other systems with quenched disorder, see, e.g., Refs. CPV-11 ; AV-11 ; PPV-09 .

### v2

# Finite-temperature topological transitions in the presence of quenched uncorrelated disorder

@@ Abstract
We address issues related to the presence of defects at finite-temperature topological transitions, in particular when defects are modeled in terms of further variables associated with a quenched disorder, corresponding to the limit in which the defect dynamics is very slow. As a paradigmatic model, we consider the classical three-dimensional lattice ${\mathbb{Z}}_{2}$ gauge model in the presence of quenched uncorrelated disorder associated with the plaquettes of the lattice, whose topological transitions are characterized by the absence of a local order parameter. We study the critical behaviors in the presence of weak disorder. We show that they belong to a new topological universality class, different from that of the lattice ${\mathbb{Z}}_{2}$ gauge models without disorder, in agreement with the Harris criterium for the relevance of uncorrelated quenched disorder when the pure system undergoes a continuous transition with positive specific-heat critical exponent.
@@ I Introduction
Statistical systems with quenched disorder, modeling the presence of defects subject to a slow dynamics, generally develops peculiar critical phenomena associated with new universality classes with respect to pure systems, such as transitions to glassy phases. Finite-temperature phase transitions and critical behaviors in disordered systems have been mostly studied within lattice spin models with various realizations of local disorder, in particular with types of disorder preserving the symmetry of the pure system, such as random-bond and random-site disorder (random-field disorder instead generally breaks the global symmetry), and spatially uncorrelated probability distributions of the quenched variables, see, e.g., Refs. Ma-book ; Nishimori-book ; Harris-74 ; EA-75 ; Nishimori-81 ; HPPV-07 ; HPPV-07-rdi ; HPPV-07-mc ; PC-99 ; Betal-00 ; KM-02 ; KR-03 ; KKY-06 ; Jorg-06 ; HPV-08 ; BFMMMY-11 ; Janus-13 ; LPP-16 ; Granato-04 ; MHT-04 ; PR-05 ; LY-07 ; VK-09 ; FMPTY-09 ; BFMS-14 ; OUK-20 ; AV-11 ; Asp-etal-16 ; Hartmann-99 ; CPV-11 ; PHP-06 ; LH-88 ; HPPV-08 ; PPV-09 ; FMMPR-16 ; CP-19 ; Nishimori-24 ; Vojta-19 .
@@
While a large amount of disorder may lead to glassy phases and glassy critical behaviors in classical three-dimensional (3D) spin systems, ferromagnetic transitions may still occur for sufficiently weak disorder. However, even though the quenched disorder does not change the symmetry-breaking pattern, the universality class of the critical behavior may differ from that of the ferromagnetic transition of the pure system. According to the Harris criterium Harris-74 , which applies to a spatially-uncorrelated quenched disorder that is effectively coupled to the energy density, ferromagnetic transitions in the presence of weak disorder may develop a new universality class when the specific-heat exponent $\alpha$ of the pure transition is positive, as it happens in 3D random-bond and random-site Ising systems Ma-book ; Nishimori-book ; HPPV-07 ; HPPV-07-rdi . On the other hand, if the pure-system continuous transition has a negative specific-heat exponent, then its universality class turns out to be stable with respect to weak disorder, as it happens in 3D random-bond O($N$) vector models for any $N\geq 2$, such as the XY and Heisenberg spin models PV-02 , corresponding to $N=2$ and $N=3$ respectively. Moreover, with increasing disorder, spin systems may develop peculiar multicritical behaviors along the so-called Nishimori line Nishimori-book ; Nishimori-81 ; LH-88 ; HPPV-07-mc ; HPPV-08 ; AV-11 , and then transitions to glassy phases, see, e.g., Refs. KKY-06 ; HPV-08 ; BFMMMY-11 ; Janus-13 ; LPP-16 ; Granato-04 ; MHT-04 ; PR-05 ; LY-07 ; VK-09 ; FMPTY-09 ; BFMS-14 ; OUK-20 ; AV-11 ; Asp-etal-16 ; CPV-11 .
@@
The effects of quenched disorder have been much less investigated at topological transitions, such as those occurring in systems with gauge symmetries, without a local order parameter driving the transition Wegner-71 ; Kogut-79 ; Sachdev-19 ; BPV-25 . A paradigmatic model for systems undergoing finite-temperature topological transitions is provided by the 3D lattice ${\mathbb{Z}}_{2}$ gauge model Wegner-71 ; Kogut-79 , whose confinement-deconfinement transition is not driven by a local order parameter but by the nonlocal order parameter associated with the area/perimeter law of the Wilson loops Wegner-71 ; Kogut-79 . 3D lattice ${\mathbb{Z}}_{2}$ gauge models have been also considered in the presence of local spatially-uncorrelated quenched disorder associated with each plaquette of the lattice DKLP-02 ; WHP-03 (denoted by RPGM, for random plaquette gauge model, in the following), i.e., with randomly chosen plaquettes having couplings with the wrong (negative instead of positive) sign. The RPGM was originally introduced in relation to the theory of the quantum error correction. However, it also provides a paradigmatic example of systems undergoing topological transitions in the presence of disorder. Some results on its temperature-disorder phase diagram were reported in Refs. WHP-03 ; OAIM-04 , but the critical behavior at the transition lines has never been analyzed. Therefore the characterization of the universality classes of its topological transitions in the presence of quenched disorder is still an open issue.
@@
In this paper we investigate the effects of disorder at finite-temperature topological transitions, focusing on the paradigmatic 3D RPGM. We recall that the pure 3D lattice ${\mathbb{Z}}_{2}$ gauge model is related to the standard 3D Ising model by a duality relation Wegner-71 ; Savit-80 , thus they share the same length-scaling critical exponent $\nu_{\cal I}\approx 0.630$ PV-02 , which implies that the specific-heat exponent is positive, $\alpha_{\cal I}\approx 0.110$. We present a numerical analysis of the critical behaviors of the 3D lattice ${\mathbb{Z}}_{2}$ gauge model for weak disorder, which is made difficult by the fact that its topological transitions do not have gauge-invariant local order parameters. We thus study the critical behaviors of this system by focusing on the scaling behavior of the gauge-invariant energy cumulants. The numerical results show that the quenched disorder destabilizes the pure-system topological transition, confirming the Harris criterium. Indeed, we observe critical behaviors belonging to another universality class with $\nu=0.82(2)$, which is significantly larger that the Ising length-scale exponent $\nu_{\cal I}\approx 0.630$ in the absence of disorder (and also different from any universality class of Ising systems with quenched disorder). In the final part of the paper we also briefly discuss generalizations of the RPGM, considering generic 3D lattice ${\mathbb{Z}}_{N}$ gauge models with weak disorder.
While in this work we focus on the effect of weak disorder at finite-temperature topological transitions, it is important to note that the presence of disorder can led to very interesting phenomenological consequences also in many quantum (i.e., zero temperature) contexts, in particular for what regards quantum critical phenomena, see, e.g., Refs. Sachdev-book-qpt ; Vojta-19 , and topological phases of matter, see, e.g., Refs. MoessnerMoore-book ; Sachdev-book-phases ; SM-09 ; GWATB-09 ; GRRF-10 ; SLJSX-12 ; GLR-13 ; TLRR-15 ; OSBS-16 ; AFL-24 ; SCMVL-24 ; ZT-26 .
@@ III Numerical methods and results
$B_{k}(K,L)\approx L^{k/\nu-3}{\cal B}_{k}(X)\left[1+L^{-\omega}{\cal B}_{k,\omega}(X)\right]+b_{k},$
@@
where the constants $b_{k}$ represent the analytic background PV-02 ; BPV-24-coH , and the $O(L^{-\omega})$ suppressed term with $\omega>0$ represents the leading scaling correction which is generally associated with the leading irrelevant renormalization group (RG) perturbation at the fixed point PV-02 (for example $\omega\approx 0.8$ at the standard Ising transition, i.e., for $q=0$). The scaling functions ${\cal B}_{k}(X)$ are universal apart from a multiplicative factor and a normalization of the argument, however they generally depend on the boundary conditions adopted.
@@
If $\omega<3/\nu-3$ then $\kappa=\omega$ and the $O(L^{-\kappa})$ term parametrizes the correction due to the leading irrelevant perturbation, see Eq. (11). If instead $\omega>3/\nu-3$ then $\kappa=3/\nu-3$, the $O(L^{-\kappa})$ term takes into account the analytic background in Eq. (11), and in this case $n_{s}=0$ should be used.
@@
instead of the definition reported in Eq. (11). The final result of this analysis are the estimates
@@ IV Conclusions
We have investigated the effects of quenched disorder at finite-temperature topological transitions, focusing on the 3D ${\mathbb{Z}}_{2}$ gauge model in the presence of an uncorrelated quenched disorder associated with the plaquettes of the lattice (RPGM). This is paradigmatic model for systems undergoing finite-temperature topological transitions Wegner-71 ; Kogut-79 ; Sachdev-19 ; BPV-25 , indeed its confinement-deconfinement transitions are not driven by any local order parameter. We report a numerical analysis of their critical behaviors for weak disorder, and for this purpose we focus on the scaling behavior of the gauge-invariant energy cumulants defined in Eq. (6). The numerical investigation has been carried out using the binomial distribution for disorder variables, but it is natural to guess that the universality class of the results is independent of the disorder distribution, as it happens for the Ising model with quenched disorder, see, e.g., Ref. HPPV-07-rdi .
@@
$\frac{1}{V}[\langle\mathcal{H}\rangle]_{w}=3-6q$ (16)
holds along the Nishimori line (5) (see also analogous behaviors Nishimori-book ; Nishimori-81 for the random-bond Ising model). Therefore, the scaling part of the cumulants vanishes along the Nishimori line, and to study the singular behavior of the free energy it is thus necessary not to move along the Nishimori line. This makes an investigation of the multicritical behavior at the Nishimori point extremely challenging from the computational point of view, since it requires the independent tuning of two different parameters (i.e., both $K$ and $q$) to approach the unknown position of the multicritical point. We also note that the critical behaviors along the deconfinement-confinement transition line below the Nishimori point, see Fig. 1, is expected to belong to a different universality class, as it occurs for other systems with quenched disorder, see, e.g., Refs. CPV-11 ; AV-11 ; PPV-09 .